(1) Explain the term electric field and line of force


(2) Derived an expression for a dipole in an electric field.


(3) Compared electrostatic and gravitational field and show difference from each other.



(4) Calculate the electronic potential energy of an electron system of an atom.The radius of the orbit of the electron.2.16*10^-11m. The charge on the electron is 1.6*10^-19C. Where 1/4Eo=9.0*10^9NM^2-2


(5) Calculate the capacity of parallel plates condenser.what will be the capacity if the space between the plates is partially filled with slap of thickness and dielectric constant Er

1
Expert's answer
2022-02-06T14:34:30-0500

1) An electric line of force is an imaginary continuous line or curve drawn in an electric field such that tangent to it at any point gives the direction of the electric force at that point. the direction of a line of force is the direction along which a small free positive charge will move along the line.

2) The electric field at a point P on axial line of dipole due to +q charge 

"E=\frac{1}{4\pi \epsilon_o}*\frac{2p}{r^3}" where p=q ,2a=dipole moment of given dipole

3) They both act between two bodies without any means of contact. However gravitational force acts on mass while the electric force acts on charge. Gravitational force are only attractive while electric field can be attractive/repulsive.

4) "r = 2.16*10^{-11}m"

"e = 1.6*10^{-19}C"

"k = 9*10^9"

"F - ?"

"W = \large k\frac{e^2}{r}=\frac{9*10^9*2.56*10^{-38}}{2.16*10^{-11}}=" "10.66*10^{-18}J"